OpenStudy (anonymous):

make p the subject of the formula q=5/3+p

OpenStudy (anonymous):

\[q=\frac{ 5 }{ 3+p }\] \[\frac{ 1 }{ q }=\frac{ 3+p }{ 5 }\] \[\frac{ 5 }{ q }-3=p\] \[p=\frac{ 5-3q }{ q }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):